# Exact Expressive Power of Transformers with Padding

### v1

@@ 2.4 Logic
We define the standard notion first-order logic over strings ($\mathsf{FO}$; cf. Merrill and Sabharwal, 2023c). $\mathsf{FO}$ formulas map strings to boolean values. Its formulas can check for token occurrences at specific string positions, and it allows quantification over positions in the input string.
@@
More formally, $\mathsf{FO}$ can be defined as follows. $\mathsf{FO}$ contains two types: indices, representing positions in the input string, and formulas, which evaluate to true or false. For an index or formula $x$, we write $\llbracket x\rrbracket^{w,v}$ for the evaluation of $x$ on string $w$ with variable assignments $v$ (a map from names to values). Indices in $\mathsf{FO}$ are integers denoting positions in the input string:
@@
It is known that the languages defined by $\mathsf{FO}$ are exactly the circuit complexity class $\mathsf{DLOGTIME}$-uniform $\mathsf{AC}^{0}$ (Mix Barrington et al., 1990). This class captures hard-attention transformers (Hao et al., 2022; Yang et al., 2024), but it is not large enough to capture soft-attention transformers (Merrill and Sabharwal, 2023a). To consider a more expressive logic capable of modeling soft-attention transformers, we can extend $\mathsf{FO}$ with majority quantifiers (Merrill and Sabharwal, 2023c). Specifically, we will define $\mathsf{FO}[\mathsf{M}^{2}]$ as $\mathsf{FO}$ extended to include a paired majority quantifier:
Definition 8 (Mix Barrington et al., 1990).
$\mathsf{FO}[\mathsf{M}^{2}]$ is $\mathsf{FO}$ extended to include the $\mathsf{M}^{2}$ quantifier, defined as follows: $\mathsf{M}^{2}(i,j).P(i,j)$ is true if $\phi(i,j)$ holds for a majority of pairs of positions $(i,j)\in[n]^{2}$ in the string.
It is known that $\mathsf{FO}[\mathsf{M}^{2}]$ defines exactly $\mathsf{DLOGTIME}$-uniform $\mathsf{TC}^{0}$ (Mix Barrington et al., 1990). It is also possible to use majority quantifiers to define addition over indices, so without loss of generality, we can assume $\mathsf{FO}[\mathsf{M}^{2}]$ formulas have no addition (unlike $\mathsf{FO}$ formulas). Moreover, it is possible to simulate $\mathsf{bit}$ in terms of $\mathsf{M}^{2}$, so we can consider $\mathsf{FO}[\mathsf{M}^{2}]$ formulas not to contain $\mathsf{bit}$, in contrast to the more standard logic $\mathsf{FO}[\mathsf{M},\mathsf{bit}]$ that also defines $\mathsf{TC}^{0}$. This makes $\mathsf{FO}[\mathsf{M}^{2}]$ a simpler target for our transformer constructions than $\mathsf{FO}[\mathsf{M},\mathsf{bit}]$.
@@ 3 Masked and Unmasked Transformers
Let $E$ be an unmasked AHAT encoder with depth $\ell\geq 1$. Then there exists a causally masked AHAT decoder $D$ with depth $\ell$ and with $\ell n$ padding tokens on input length $n$ that is equivalent to $E$ in the following sense: the final $n$ outputs of $D$ match the original $n$ outputs of $E$.
@@
The idea is that $D$ will unroll a sequence of $\ell$ “blocks” on the padding length dimension, each of width $n$. Each block will attend to the previous block in the previous layer and can thus see all tokens despite causal masking.
To implement the block construction, we first compute $\phi(n)$. We then use this to compute $\phi(b_{i})$, where $b_{i}=\lfloor i/n\rfloor\in[1,\ell]$, which represents the block that each token belongs to. We also compute $\phi(b_{i}-1)\in[0,\ell-1]$ using a head that ignores the bos token. Next, we modify each original head from $U$ to have an additional term in the query/key product. The query is $C\phi(b_{i}-1)$ and the key is $\phi(b_{j})$, where $C$ is a large weight. We set $C$ to a fixed large value (independent of $n$) such that this term dominates all other terms in the inner product computation at each token. As a result, the head is constrained over keys in the context where $b_{j}=b_{i}-1$, i.e., the keys from the last block. Within these keys, it computes exactly the same AHAT output as the original masked unmasked head. In this way, we are able to simulate a causally masked transformer with an unmasked transformer. ∎
This allows us to relate various masking variants of $\mathsf{AHAT}$ classes as follows:
@@
Unmasked and mixed-masked padded transformers (i.e., uAHATs and mAHATs) can be simulated by causally masked transformers (AHATs). Specifically, for any $d,k\in\mathbb{N}$, the following holds for the corresponding problem classes:
@@
We defer the proof of this to Appendix A.
4 Fixed-Depth Padded Transformers Recognize Exactly $\mathsf{TC}^{0}$
@@
An $\mathsf{FO}[\mathsf{M}^{2}]$ formula with $k\geq 1$ distinct variables and nesting depth $\ell$ can be computed in $\mathsf{uAHAT}^{0}_{k}$ (and hence in $\mathsf{AHAT}^{0}_{k}$) with (fixed) depth $\ell$.
@@
We will store all possible configurations of $k$ variables using $n^{k}$ padding tokens. Each token corresponds to a specific assignment of all the variables, which we denote $v$ (cf. Definition 8). We will present an inductive transformer construction that uses a single layer to compute the boolean variable of each formula and the integer value of each numerical expression in $\mathsf{FO}[\mathsf{M}^{2}]$, assuming the constituent formulas and values were already computed at the previous layer. Since each $\mathsf{FO}[\mathsf{M}^{2}]$ formula consists of a finite number of subformulas and values, we can accumulate all constituents in the residual stream in order to compute any larger formula.
In more detail, let $\llbracket x\rrbracket^{w,v}$ be the value of a formula or numerical value $x$ on string $w$ under assignment $v$. We identify each padding token $v$ with a variable assignment $v$ so that, for every $P$, token $v$ stores $\llbracket P\rrbracket^{w,v}$ in the residual stream with a scalar whose sign indicates a truth value. For a numerical value $i$, token $v$ will represent $\llbracket i\rrbracket^{w,v}$ as a small vector $\phi(\llbracket i\rrbracket^{w,v})$. where $\phi$ is the layer-norm hash (Definition 3). We show how to evaluate each constituent of an $\mathsf{FO}[\mathsf{M}^{2}]$ formula. As mentioned after Definition 8, addition over indices and $\mathsf{bit}$ are subsumed by $\mathsf{M}^{2}$, so we do not need to simulate them.
@@
    Variables. At each token $v$, we will compute $\phi(\llbracket i\rrbracket^{w,v})$, a representation of the value of variable $i$ under assignment $v$. We view the integer $v$ as a tuple of $k$ values in $[1,n]$, so we will take a “projection” to retrieve $v[i]$ using quotient and remainder operations (Merrill and Sabharwal, 2024, Lemma 3.1). More formally, if $i$ is the $m$-th variable, compute
    $\phi(\lfloor v/n^{m-1}\rfloor\bmod n)=\phi(\llbracket i\rrbracket^{w,v}).$
    Thus, we conclude that we can compute $\phi(\llbracket i\rrbracket^{w,v})$ at each token $v$.
@@
    Constants. To compute the constant $1$ or $n$ at each configuration $v$, we can attend from each $v$ to the first and last input token and retrieve its position to get $\phi(1)$ or $\phi(\lvert w\rvert)$.
@@
    Standard Quantifiers. Assume we have $\llbracket P\rrbracket^{w,v}$ stored at each configuration $v$. Then, at each $m$, we want to resolve the quantifier $\mathsf{Q}$ over the set $\{\llbracket P\rrbracket^{w,v|i=m}\}_{m=1}^{n}$, where $v|i=m$ denotes $v$ with $i$ overriden to have value $m$. We will count $c$, the number of $m$ such that $P^{v|i=m}$ holds. More formally, let $j_{1},\ldots,j_{k-1}$ be the set of variables excluding $i$. Then we can attend from $v$ over $v^{\prime}$ with query $\langle\phi(j_{1}^{v},\ldots,\phi(j_{k-1}^{v}\rangle$, key $\langle\phi(j_{1}^{v^{\prime}},\ldots,\phi(j_{k-1}^{v^{\prime}})\rangle$, and value $P^{v^{\prime}}$ to retrieve $c/n$. Finally, we threshold $c/n$ against $\frac{1}{2n}$ (for $\exists$) or $\frac{2n-1}{2n}$ (for $\forall$) to resolve $(\mathsf{Q}i.\phi)^{v}$.
@@
    Paired Majority Quantifiers. Assume we have $P^{v}$ stored at each configuration $v$. We want to compute $(\mathsf{M}(i,j).P)^{v}$ for any two variables $i,j$ already represented. The idea slightly generalizes the construction for single-variable quantifiers: we will use attention to count $c$, the number of assignments where $P^{v|i=m,j=\ell}$ is true. Formally, let $j_{1},\ldots,j_{k-2}$ be the set of variables excluding $i,j$. Then we can attend from $v$ over $v^{\prime}$ with query $\langle\phi(j_{1}^{v},\ldots,\phi(j_{k-2}^{v}\rangle$, key $\langle\phi(j_{1}^{v^{\prime}},\ldots,\phi(j_{k-2}^{v^{\prime}})\rangle$, and value $P^{v^{\prime}}$ to retrieve $c/n^{2}$. Finally, we threshold $c/n^{2}$ against $\frac{1}{2}$ to resolve $(\mathsf{M}^{2}(i,j).P)^{v}$.
In conclusion, we can compute any $\mathsf{FO}[\mathsf{M}^{2}]$ formula in $\mathsf{uAHAT}^{0}_{k}$ by inductively computing its constituent formulas and storing their values over $k$ variables using $n^{k}$ padding tokens. The result extends to $\mathsf{AHAT}^{0}_{k}$ by applying Proposition 1. ∎
Our construction for Lemma 2 somewhat resembles (though is distinct from) a result of Lange (2004, Corollary 6.8) that any problem in $\mathsf{TC}^{0}$ can be reduced to majority logic via a transformation that appends some extra tokens to the input. However, their transformation is not padding since it appends some “non-blank” tokens, and their result also does not clearly apply to transformers.
Combined with previous results about transformers being in $\mathsf{FO}$-uniform $\mathsf{TC}^{0}$, Lemma 2 yields an exact characterization of constant-depth transformers with padding:
@@
It is known that $\mathsf{AHAT}^{0}\subseteq\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{0}$ (Merrill and Sabharwal, 2023a, c; Chiang, 2025), and this generalizes to transformers with padding tokens (Pfau et al., 2024). We will show $\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{0}\subseteq\mathsf{uAHAT}^{0}_{*}$. Mix Barrington et al. (1990, Proposition 10.3) proved that $\mathsf{FO}$-uniform $\mathsf{TC}^{0}$ is definable by $\mathsf{FO}$ formulas with $\mathsf{M}^{2}$ quantifiers: notably, the $\mathsf{bit}$ predicate is not necessary when using $\mathsf{M}^{2}$ quantifiers. Thus, Lemma 2 establishes that $\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{0}\subseteq\mathsf{uAHAT}^{0}_{*}$. Finally, from Proposition 1, we have $\mathsf{uAHAT}^{0}_{*}\subseteq\mathsf{AHAT}^{0}_{*}$. Hence, $\mathsf{uAHAT}^{0}_{*}=\mathsf{AHAT}^{0}_{*}=\mathsf{FO}$-uniform $\mathsf{TC}^{0}$. ∎
A technical hurdle to obtaining this characterization in prior work was simulating the $\mathsf{bit}$ predicate in standard definitions of $\mathsf{TC}^{0}$ (Pfau et al., 2024). Our results circumvent this by instead relating transformers to $\mathsf{FO}[\mathsf{M}^{2}]$, which is equivalent to standard $\mathsf{FO}[\mathsf{M},\mathsf{bit}]$ (Mix Barrington et al., 1990). It thus follows from Theorem 1 that padded transformers can also simulate $\mathsf{bit}$ as well as all of $\mathsf{FO}$, which will be useful in the following section for simulating $\mathsf{FO}$ reductions.
5 Logd-Looped Padded Transformers Recognize Exactly $\mathsf{TC}^{d}$

### v2

@@ 2.4 Logic
We define the standard notion first-order logic over strings ($\mathsf{FO}$; cf. merrill2023logic). $\mathsf{FO}$ formulas map strings to boolean values. Its formulas can check for token occurrences at specific string positions, and it allows quantification over positions in the input string. More formally:
@@
$\mathsf{FO}$ contains two types: indices, representing positions in the input string, and formulas, which evaluate to true or false. For an index or formula $x$, we write $\llbracket x\rrbracket^{w,v}$ for the evaluation of $x$ on string $w$ with variable assignments $v$ (a map from names to values). Indices in $\mathsf{FO}$ are integers denoting positions in the input string:
@@
It is known that the class of languages defined by $\mathsf{FO}$ is exactly the circuit complexity class $\mathsf{DLOGTIME}\textrm{-uniform}\ \mathsf{AC}^{0}=\mathsf{FO}\textrm{-uniform}\ \mathsf{AC}^{0}$ (barrington-1990-uniformity). This class captures the languages recognized by unique hard-attention transformers (hao-etal-2022-formal; yang2024masked), but it is not large enough to capture soft-attention transformers (merrill-sabharwal-2023-parallelism). To consider a more expressive logic capable of modeling soft-attention transformers, we can extend $\mathsf{FO}$ with majority quantifiers (merrill2023logic). Specifically, we will define $\mathsf{FO}+\mathsf{M}^{2}$ as $\mathsf{FO}$ extended to include a paired majority quantifier:
Definition 8 (barrington-1990-uniformity).
$\mathsf{FO}+\mathsf{M}^{2}$ is $\mathsf{FO}$ extended to include the $\mathsf{M}^{2}$ quantifier, defined as follows: $\mathsf{M}^{2}(i,j).P(i,j)$ is true if $\phi(i,j)$ holds for a majority of pairs of positions $(i,j)\in[n]^{2}$ in the string.
It is known that $\mathsf{FO}+\mathsf{M}^{2}$ defines exactly $\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{0}$ (barrington-1990-uniformity, Theorem 10.2, which uses paired majority). It is also possible to use majority quantifiers to define addition over indices, so without loss of generality, we can assume $\mathsf{FO}+\mathsf{M}^{2}$ formulas have no addition (unlike $\mathsf{FO}$ formulas). Moreover, it is possible to simulate $\mathsf{bit}$ in terms of $\mathsf{M}^{2}$, so we can consider $\mathsf{FO}+\mathsf{M}^{2}$ formulas not to contain $\mathsf{bit}$, in contrast to the more standard logic $\mathsf{FO}+\mathsf{M}[\mathsf{bit}]$ that also defines $\mathsf{TC}^{0}$. This makes $\mathsf{FO}+\mathsf{M}^{2}$ a simpler target for our transformer constructions than $\mathsf{FO}+\mathsf{M}[\mathsf{bit}]$.
@@ 3 Masked and Unmasked Transformers
Let $E$ be an unmasked (with position encoding $1/i$) AHAT encoder with depth $\ell\geq 1$. Then there exists a causally masked AHAT decoder $D$ (without any position encoding) with depth $\ell$ and with $\ell n$ padding tokens on input length $n$ that is equivalent to $E$ in the following sense: the final $n$ outputs of $D$ match the original $n$ outputs of $E$.
@@
We first observe that the causally masked decoder $D$ can compute $1/i$ (the position encoding used in the unmasked encoder $E$) by attending uniformly with value $1$ only for the beginning-of-sequence symbol. To simulate unmasked attention with causally masked attention via padding, the idea is for $D$ to unroll a sequence of $\ell$ “blocks” (one for each of the $\ell$ layers of $E$) along the padding length dimension, each of width $n$. Each block will attend to the previous block in the previous layer and can thus see all tokens despite causal masking.
[…]
Since causally masked heads of a mixed-masked transformer can be trivially simulated by a causally masked transformer, Lemma˜1 allows us to relate various masking variants (proof in Appendix˜C):
@@
Unmasked (with position encoding $1/i$ or $i/n$; cf. Lemma˜10 in §C) and mixed-masked padded transformers (i.e., uAHATs and mAHATs) can be simulated by causally masked transformers (AHATs). Specifically, for any $d,k\in\mathbb{N}$, the following holds for the corresponding problem classes:
@@
4 Fixed-Depth Padded Transformers Recognize Exactly $\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{0}$
@@
An $\mathsf{FO}+\mathsf{M}^{2}$ formula with $k\geq 1$ distinct variables and nesting depth $\ell$ can be computed in $\mathsf{uAHAT}^{0}_{k}$ (and hence in $\mathsf{AHAT}^{0}_{k}$) with (fixed) depth $\ell$.
@@
We will store all $n^{k}$ possible configurations of $k$ variables of the formula using $n^{k}$ padding tokens, where each token corresponds to a specific configuration of all the variables, which we denote $v$ (cf. Definition˜7). We will present an inductive transformer construction that uses a single layer to compute the boolean variable of each formula and the integer value of each numerical expression in $\mathsf{FO}+\mathsf{M}^{2}$, assuming the constituent formulas and values were already computed at the previous layer. Since each $\mathsf{FO}+\mathsf{M}^{2}$ formula consists of a fixed number of subformulas and values, we can accumulate all constituents in the residual stream in order to compute any larger formula.
In more detail, let $\llbracket x\rrbracket^{w,v}$ be the value of a formula or numerical value $x$ on string $w$ under assignment $v$, with $k$ total variables. We will identify each variable assignment with an integer $v\in[n^{k}]$, so that, for every $P$, padding token $v$ stores $\llbracket P\rrbracket^{w,v}$ in the residual stream with a scalar whose sign indicates a truth value. For a numerical value $i$, token $v$ will represent $\llbracket i\rrbracket^{w,v}$ as a small vector $\phi(\llbracket i\rrbracket^{w,v})$. where $\phi$ is the layer-norm hash (Definition˜3). We show how to evaluate each constituent of an $\mathsf{FO}+\mathsf{M}^{2}$ formula. As mentioned after Definition˜8, addition over indices and $\mathsf{bit}$ are subsumed by $\mathsf{M}^{2}$, so we do not need to simulate them.
@@
    Constants. To compute the constant $1$ or $n$ at each configuration $v$, we can attend from each $v$ to the first and last input token and retrieve its position to get $\phi(1)$ or $\phi(n)$, where $n=\lvert w\rvert$.
@@
    Variables. At each token $v$, we will compute $\phi(\llbracket i\rrbracket^{w,v})$, a representation of the value of variable $i$ under assignment $v$. We view the integer $v\in[n^{k}]$ as a tuple of $k$ values $v[1],\ldots,v[k]\in[n]$. We have that $\llbracket i\rrbracket^{w,v}=v[i]$. To get this, we first compute $\phi(v)$ and then take a “projection” to retrieve $\phi(v[i])$ using quotient and remainder operations (merrill-2025-little, Lemma 3.1). More formally, we compute $\phi(n)$ where $n=\lvert w\rvert$, divide i1i-1 times by $n$, and then take the remainder $\bmod\ n$ to obtain:
    $\phi(\lfloor v/n^{i-1}\rfloor\bmod n)=\phi(v[i]).$
    Thus, we conclude that we can compute $\phi(v[i])=\phi(\llbracket i\rrbracket^{w,v})$ at each token $v$.
@@
    Standard Quantifiers. Assume we have $\llbracket P\rrbracket^{w,v}$ stored at each configuration $v$. Then, at each $m$, we want to resolve the quantifier $\mathsf{Q}$ over the set $\{\llbracket P\rrbracket^{w,v|i=m}\}_{m=1}^{n}$, where $v|i=m$ denotes $v$ with $i$ overriden to have value $m$. We will count $c$, the number of $m$ such that $P^{v|i=m}$ holds. More formally, let $j_{1},\ldots,j_{k-1}$ be the set of variables excluding $i$. Then we can attend from $v$ over $v^{\prime}$ with query $\langle\phi(j_{1}^{v}),\ldots,\phi(j_{k-1}^{v})\rangle$, key $\langle\phi(j_{1}^{v^{\prime}}),\ldots,\phi(j_{k-1}^{v^{\prime}})\rangle$, and value $P^{v^{\prime}}$ to retrieve $c/n$. Finally, we threshold $c/n$ against $\frac{1}{2n}$ (for $\exists$) or $\frac{2n-1}{2n}$ (for $\forall$) to resolve $\llbracket\mathsf{Q}i.\phi\rrbracket^{w,v}$.
@@
    Paired Majority Quantifiers. Assume we have $\llbracket P\rrbracket^{w,v}$ stored at each configuration $v$. We want to compute $\llbracket\mathsf{M}(i,j).P\rrbracket^{w,v}$ for any two variables $i,j$ already represented. The idea slightly generalizes the construction for single-variable quantifiers: we will use attention to count $c$, the number of assignments where $\llbracket P\rrbracket^{w,v|i=m,j=\ell}$ is true. Formally, let $j_{1},\ldots,j_{k-2}$ be the set of variables excluding $i,j$. Then we can attend from $v$ over $v^{\prime}$ with query $\langle\phi(j_{1}^{v}),\ldots,\phi(j_{k-2}^{v})\rangle$, key $\langle\phi(j_{1}^{v^{\prime}}),\ldots,\phi(j_{k-2}^{v^{\prime}})\rangle$, and value $\llbracket P\rrbracket^{w,v^{\prime}}$ to retrieve $c/n^{2}$. Finally, we threshold $c/n^{2}$ against $\frac{1}{2}$ to resolve $\llbracket\mathsf{M}^{2}(i,j).P\rrbracket^{w,v}$.
In conclusion, we can compute any $\mathsf{FO}+\mathsf{M}^{2}$ formula in $\mathsf{uAHAT}^{0}_{k}$ by inductively computing its constituent formulas and storing their values over $k$ variables using $n^{k}$ padding tokens. The result extends to $\mathsf{AHAT}^{0}_{k}$ by applying Proposition˜1. ∎
Our construction for Lemma˜2 somewhat resembles (though is distinct from) a result of lange-2004-results that any problem in $\mathsf{TC}^{0}$ can be reduced to majority logic via a transformation that appends some extra tokens to the input. However, their transformation is not padding since it appends some “non-blank” tokens, and their result also does not clearly apply to transformers.
Combined with previous results about transformers being in $\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{0}$, Lemma˜2 yields an exact characterization of constant-depth transformers with padding:
@@
It is known that $\mathsf{AHAT}^{0}\subseteq\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{0}$ (merrill-sabharwal-2023-parallelism; merrill2023logic; chiang2025transformers), and this generalizes to transformers with padding tokens (pfau-2024-think). We will show $\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{0}\subseteq\mathsf{uAHAT}^{0}_{*}$. barrington-1990-uniformity proved that $\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{0}$ is definable by $\mathsf{FO}$ formulas with $\mathsf{M}^{2}$ quantifiers: notably, the $\mathsf{bit}$ predicate is not necessary when using $\mathsf{M}^{2}$ quantifiers. Thus, Lemma˜2 establishes that $\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{0}\subseteq\mathsf{uAHAT}^{0}_{*}$. Finally, from Proposition˜1, we have $\mathsf{uAHAT}^{0}_{*}\subseteq\mathsf{AHAT}^{0}_{*}$. Hence, $\mathsf{uAHAT}^{0}_{*}=\mathsf{AHAT}^{0}_{*}=\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{0}$. ∎
A technical hurdle to obtaining this characterization in prior work was simulating the $\mathsf{bit}$ predicate in standard definitions of $\mathsf{TC}^{0}$ (pfau-2024-think). Our results circumvent this by instead relating transformers to $\mathsf{FO}+\mathsf{M}^{2}$, which is equivalent to standard $\mathsf{FO}[\mathsf{M},\mathsf{bit}]$ (barrington-1990-uniformity). It thus follows from Theorem˜1 that padded transformers can also simulate $\mathsf{bit}$ as well as all of $\mathsf{FO}$, which will be useful in the following section for simulating $\mathsf{FO}$ reductions.
Lastly, we note that the arguments leading to Theorem˜1 work for both logarithmic and polynomial precision (cf. Appendix˜A) transformers. This shows that these two levels of precision lead to identical power for fixed-depth padded transformers.
5 Logd-Looped Padded Transformers Recognize Exactly $\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{d}$